OpenStudy (anonymous):

-5x*+6y*+2 x=0 , y=3

OpenStudy (anonymous):

326

jhonyy9 (jhonyy9):

0+18+2=20

OpenStudy (anonymous):

agree with johnyy9, -5(0)+6(3)+2 = 0+18+2=20 only thing is dont know what asterisk after x and y mean?